# Xpikeformer: Hybrid Analog-Digital Hardware Acceleration for Spiking Transformers

### v1

@@ III-A ANN Transformer Accelerators
III-A1 CMOS-based Accelerators
The development of CMOS-based ANN transformer accelerators primarily aims to enhance computational efficiency. Various strategies have been explored in the literature to achieve this goal. Mixed-precision techniques [30], algorithm-specific quantization [31], and adaptive quantization methods based on input data characteristics [32] have all been utilized. One approach decomposes major operations into a single dot product primitive, allowing for unified and efficient execution, thereby significantly improving throughput [33]. Additionally, an optimized datapath specifically designed for attention mechanisms has been developed, reducing latency and enhancing throughput [32]. Another method addresses inefficiencies in the decoding process by skipping redundant computations and adopting a sparse matrix format to maintain high utilization of MACs [34]. Despite these advancements, the current solutions to the problem of data transfer overhead remain inadequate.
@@ III-A1 CMOS-based Accelerators
III-A2 AMIC-Integrated Accelerators
Advancements in AIMC have significantly enhanced the efficiency and performance of accelerators for ANN transformer models. The most straightforward method leverages AIMC for efficient MVMs in the feedforward layers while retaining the MHSA in general-purpose computing units [35]. This hybrid approach, although shown to achieve software-equivalent accuracy on NLP tasks, still experiences certain energy inefficiencies due to the high-precision digital computation required for MHSA [36].
Efforts have been made to implement MHSA with AIMC to accelerate the dot-product attention, which necessitates online programming to write the intermediate results $K$ and $V$ into crossbars [19]. Despite a dedicated sub-matrix pipeline designed by the authors, this programming process introduces significant latency and power consumption. Additionally, the wear and degradation of NVM devices render the online-programming approach impractical. A fully circuit-based implementation proposes using functional circuit modules for dot-product attention, softmax, normalization, and activation functions [20], aligning with the feedforward implementations in the crossbar. However, analog circuits introduce significant non-linearity and noise, and require even more power to operate compared to digital complementary metal-oxide-semiconductor (CMOS)-based computations.
Alternative architectures integrate near-memory computing (NMC) or static random-access memory (SRAM)-based digital in-memory computing (DIMC) to manage MHSA workloads [37, 38]. A hybrid in-memory accelerator utilizes eight-transistor (8T)-SRAM to compute MHSA with reduced leakage, employing a sequence-blocking dataflow to enhance hardware utilization and reduce execution time [21]. However, these schemes still require the key (or query) and value matrices to be written to the SRAM array during inference, causing extra latency and energy consumption. Moreover, heavy computing units are still needed to compute the softmax function.
III-B Spiking Transformer Accelerators
Despite the extensive research on hardware accelerators for SNNs, particularly for LIF models [39, 40, 41], the advancement in accelerating spiking transformers remains algorithmic and software-focused. One pioneering study in this domain proposes a spiking MHSA that captures sparse visual features using spike-form $Q$, $K$, and $V$ without employing softmax, achieving accuracy comparable to ANNs in image classification tasks [10, 12]. Another significant contribution [11] presents a hardware-friendly spike-driven residual learning architecture designed to eliminate the non-spike computations introduced by residual connections. Building on these advancements, [16] emphasizes sparse addition operations instead of multiplication in MHSA, further reducing computational energy. However, these aforementioned studies, being at the software and algorithmic level, can be inefficient when deployed on general-purpose computing platforms like CPUs and GPUs due to their intrinsic use of high-precision resources [28]. Moreover, none of these studies have addressed the data transfer overhead, leaving an important aspect of efficiency improvement unexplored. Although FPGA-based spiking attention accelerators offer a potential solution, this area remains underexplored.
IV Xpikeformer Architecture
Xpikeformer is a hybrid analog-digital hardware architecture specifically designed for spiking transformers. In the Xpikeformer architecture, two different types of computing engines are employed to handle different operations within spiking transformers, leveraging their respective strengths. The feed-forward, embedding, and fully connected layers are implemented within the AIMC engine, which primarily consists of spiking neuron tiles that include multiple memristor crossbars for storing synaptic weights. The peripheral circuits are optimized to perform accumulation and LIF functionalities without intermediate buffering. The MHSA mechanism is executed by the SSA engine, a framework that leverages stochastic computing to perform dot-product attention operations in the spiking domain. The SSA engine is a purely CMOS-based architecture that uses logical AND gates instead of resource-intensive arithmetic units.
@@ IV-A AIMC Engine
The AIMC engine is composed of multiple spiking neuron tiles. These tiles are virtually grouped to ensure that each group has sufficient memory to store the weights for one feedforward or fully connected layer. Each spiking neuron tile includes an input and an output buffer, multiple SAs, and a group of LIF units. A shared DAC unit is deployed on each spiking neuron tile to facilitate the initial programming of the SAs, allowing each device to be individually programmed sequentially to achieve multi-bit precision on a single memristor. This DAC unit is bypassed during inference.
IV-A1 Memristor Crossbars
Each SA within the spiking neuron tile primarily consists of a fixed-sized memristor crossbar, along with multiplexed readout units, as introduced in Section II-D. We impose a restriction on the size of crossbar arrays because excessively large sizes result in increased interconnect resistance and capacitance, higher power consumption, signal degradation, and increased complexity of control circuitry. Consequently, the weights of a feedforward or fully connected layer are most likely distributed across multiple SAs or even several spiking neuron tiles.
We choose the phase change memory (PCM) device to implement the crossbars in SAs due to its multi-level cell capability (up to 4 bits per device), high endurance, fast read and write speed, and compatibility with CMOS technology [42, 43, 44]. Detailed SAs configuration parameters of this work are shown in Table II. Nevertheless, as an analog device, PCM suffers from thermal noise, $1/f$ noise, device variability, and quantization effects. A significant challenge associated with PCM is conductance drift, where the conductance of the memory cell changes over time after it has been programmed. This drift can lead to inaccuracies in stored weights and degrade the performance of the neural network, particularly in terms of precision and reliability of the computations. Methods to address the effects caused by these nonidealities will be introduced in Section V.
TABLE II: Configuration of Synaptic Arrays
@@ IV-A2 Weight Mapping and Tile Architecture
To efficiently manage this distribution and avoid the overhead of storing non-binary pre-activations, we adopt a row-block-wise mapping strategy. For example, with a crossbar size of $128\times 128$ cells and a weight matrix of $384\times 512$, the weight matrix must be divided into 12 $128\times 128$ submatrices. As shown in Fig. 4, the row-block-wise mapping strategy ensures that the same row of submatrices is mapped to the SAs within a single spiking neuron tile. For instance, the 129 to 256 rows of the weight matrix are mapped to four SAs (2-1 to 2-4) within one spiking neuron tile. The input $512\times 1$ vector is then split into four $128\times 1$ subvectors and fed into SA 2-1 to SA 2-4 respectively. At each column output of an SA, only a local sum of the vector dot product can be obtained. To avoid storing this local sum, the outputs of the same column in SA 2-1 to SA 2-4 are routed to the same LIF unit. Thus, all local sums can be directly accumulated within the LIF unit using a carry-save adder (CSA) to get one output feature, i.e. pre-activation. The pre-activation is then added to the current membrane potential via an adder. The new membrane potential overwrites the previous one in a shift register. This value is compared to a threshold stored in the threshold register using a comparator. If the membrane potential exceeds the threshold, the comparator generates a ‘1’ and resets the shift register. For each time step, the shift register is right-shifted by one bit, representing a leak factor of $\beta=0.5$. The output from the comparator is taken as the output of a spiking neuron, sent to the output buffer, and then sent back to the shared SRAM.
Figure 4: A illustration of the row-block-wise mapping strategy of synaptic weights.
The row-block-wise mapping strategy is compatible with ADC sharing in SAs. In Fig. 4, the sharing ratio is 8, which means each SA has $128/8=16$ readout units. Thus, SA 2-1 to SA 2-4 are connected with 16 LIF units. The address decoding strategy of the MUX decoder is set as identical over SAs to ensure alignment between local sums. At each MUX decoding cycle, 16 of the 128 output features can be obtained.
By implementing the row-block-wise mapping strategy and utilizing the designed LIF units within the spiking neuron tiles, this design eliminates the significant overhead associated with storing and reading non-binary pre-activations repeatedly, resulting in substantial energy efficiency.
@@ IV-B SSA Engine
Stochastic spiking attention (SSA) is a software-hardware co-design of the spike-based attention mechanism aimed at minimizing the computing and memory access overhead. SSA is proposed based on the observation that the binary, stochastic nature of the queries $Q$, keys $K$, and values $V$ that are projected from Bernoulli-encoded embeddings support efficient multiplication using stochastic computing. We introduce the SSA engine following a bottom-to-top procedure.
@@ IV-B1 SSA Algorithm
where LIF($\cdot$) represents the per-entry application of LIF neuron as described in Section II-C2. While LIF neurons are stateful, the proposed SSA replaces the LIF neurons with the Bernoulli neuron layer (BNL), computing the attention results as:
@@
A detailed description of the SSA algorithm is shown in Algorithm 1. Owing to the binary nature of the representations $Q^{t}$, $K^{t}$, and $V^{t}$, the matrix multiplications in (6) are realized by lightweight logic AND operations and additions instead of real multiplications. In step 5, the $N\times N$ binary attention scores $S^{t}$ are obtained by first calculating the dot-product between $Q^{t}$ and $K^{t}$. This is done by summing the results of logical AND operations across the dimension $d_{K}$. The result is then normalized and used for Bernoulli encoding. Step 9 applies a similar logical AND to compute the matrix multiplication between $S^{t}$ and $V^{t}$ and generate probabilistically encoded attention result $A^{t}$.
@@
9:          $A^{t}_{d,n}\sim\mathrm{Bern}\left(\frac{1}{N}\sum_{n^{\prime}=1}^{N}A^{t}_{n,n^{\prime}}\land V^{t}_{d,n^{\prime}}\right)$
@@ IV-B2 SSA Tile
Each SSA tile is composed of $N^{2}$ physical stochastic attention cells (SACs) arranged as an $N\times N$ array. The $N^{2}$ parallelization level in SACs offers significant benefits. First, it enables simultaneously sending each query to each key-value pair. This is achieved by simultaneously streaming $K^{t}$ and $V^{t}$ across columns and $Q^{t}$ across rows via 1-bit data buses. Furthermore, it allows for the parallel computation of all $N\times N$ elements of the attention score $S^{t}$ matrix. The $(i,j)$-th element of matrix $S^{t}$ is computed within the corresponding $(i,j)$-th SAC. This efficient data streaming through the SSA block eliminates the need for writing/reading intermediate data from the memory. Since SOTA transformers for edge AI applications have $N$ in the range of $16-128$, we have employed $N^{2}$ physical SAU units in our design.
@@
To calculate the attention output $A^{t}$, the generated attention scores $S^{t}_{i,j}$ are held in the SAC for $d_{K}$ clock cycles. The local attention result between $S^{t}_{i,j}$ and the $j$-th row of $V^{t}$ is calculated using another AND gate, serving as the output of the SAC. To avoid using external delays for $V^{t}$, a $d_{K}$-bit shift register operating on a first-in-first-out basis is deployed in each SAC to temporarily buffer $V^{t}$ and align it with $S^{t}$. This allows for the simultaneous streaming of $Q^{t}$, $K^{t}$, and $V^{t}$ within the SSA block, facilitating pipelining over time steps. The summation of the local attention results is achieved by adding the outputs of each column of SACs using an $N$-input binary adder. The sum is then sent to another Bernoulli encoder, where it is normalized by $N$ and used as a probability to generate the Bernoulli samples. The $i$-th row of $A^{t}$ is generated sequentially by holding $S^{t}_{i,j}$ in the $(i.j)$-th SAU when streaming the $j$-th row of $V^{t}$. As a result of this procedure, the $i$th column within the SSA tile sequentially outputs the $i$-th row of the attention result $A^{t}$. With $N$ rows of SACs operating in parallel, the entire matrix $A^{t}$ is calculated column by column by one SSA tile.
@@
The Bernoulli encoder in our design uses a comparator to compare the input integer with a pseudo-random number (PRN) generated from an external source. Notably, we do not perform normalization in hardware; instead, we directly compare the unnormalized input integer $I$ with a random integer sampled uniformly from $(0,I_{\text{max}}]$. In our design, $I_{\text{max}}$ is set to $d_{K}$ within SACs and to $N$ outside of SACs. Choosing $d_{K}$ and $N$ as powers of two can streamline the hardware design by allowing us to directly use a fixed-point random integer sampled from a uniform distribution for the comparison. This approach simplifies the design and enables the use of power-efficient Linear Feedback Shift Register (LFSR)-based PRN generators.
@@ IV-B3 SSA Engine
The SSA engine is designed for efficient execution of self-attention mechanisms in transformer models. It consists of multiple SSA tiles, each serving as an individual attention head. These tiles are stateless, allowing them to be reused across different layers of the Transformer, which significantly enhances both area and energy efficiency. Additionally, the SSA engine incorporates an LFSR array that generates all the necessary PRNs for the Bernoulli neurons, ensuring a consistent supply of random numbers required for the stochastic elements of the model.
To further optimize power and area efficiency, the SSA engine implements a custom reuse strategy for random number generation, as described in [45]. Here we implement a 32-bit LFSR and tap all 4 bytes from the LFSR instead of just the lowest 8-bits [46]. This approach maximizes the utilization of generated random numbers, conserving both power and silicon area. By leveraging these design strategies, the SSA engine achieves a balance of high performance and resource efficiency.

### v2

@@ III-A ANN Transformer Accelerators
III-A1 Digital Accelerators
The development of digital (fully CMOS-based) accelerators for ANN-based transformers primarily focuses on enhancing computational efficiency. Various strategies have been explored to achieve this, including mixed-precision techniques [33], algorithm-specific quantization [34], and adaptive quantization methods based on input data characteristics [35]. One approach decomposes major operations into a single dot product primitive, enabling unified and efficient execution, significantly improving throughput [36]. Another optimization involves a dedicated data path for attention mechanisms, reducing latency and further enhancing throughput [35]. Additionally, an approach to decoding inefficiencies skips redundant computations and adopts a sparse matrix format, maximizing MAC utilization [37]. Despite these advancements, existing solutions remain inadequate in addressing data transfer overhead, which continues to be a major bottleneck.
III-A2 AMIC-Integrated Accelerators
Advancements in AIMC have enhanced efficiency and performance in ANN transformer accelerators. A hybrid approach employs AIMC for feedforward layers while executing MHSA on general-purpose units, achieving software-equivalent accuracy for NLP tasks but retaining energy inefficiencies due to high-precision digital computations [38, 39].
Efforts to implement MHSA with AIMC require online programming to store intermediate results ($K$ and $V$) into crossbars, introducing significant latency, power consumption, and impracticality due to NVM wear [22]. Circuit-based implementations replace online programming with analog modules for dot-product attention, softmax, and activation functions, aligning with crossbar feedforward layers. However, analog circuits suffer from non-linearity, noise, and higher power demands compared to digital designs [23].
Alternative architectures explore near-memory computing (NMC) or digital in-memory computing (DIMC) using SRAM for MHSA workloads. An 8T-SRAM-based hybrid accelerator reduces leakage and employs sequence-blocking dataflows to enhance hardware utilization and reduce execution time [40, 41, 24]. However, these designs still require writing key and value matrices to SRAM during inference, adding latency and energy overhead, while also relying on compute-intensive softmax operations.
III-B Advances Towards Efficient Spiking Transformers
Despite extensive research on hardware accelerators for SNNs, particularly for LIF models [42, 43, 44], efforts to accelerate spiking transformers have remained largely algorithmic and software-focused. One pioneering study in this domain proposes a spiking MHSA that captures sparse visual features using spike-form $Q$, $K$, and $V$ without employing softmax, achieving ANN-comparable accuracy in image classification tasks [13, 15]. Another significant contribution presents a hardware-friendly spike-driven residual learning architecture that eliminates non-spike computations from residual connections [14]. Building on these advancements, [19] further reduces computational energy by replacing multiplications with sparse addition operations in MHSA. However, these studies remain at the software and algorithmic level, making them inefficient on general-purpose computing platforms like CPUs and GPUs [31]. Furthermore, none have addressed the data transfer overhead, leaving a critical aspect of efficiency improvement unexplored.
IV Xpikeformer Architecture
Xpikeformer is a hybrid analog-digital architecture designed for spiking transformers. As shown in Fig. 3, Xpikeformer integrates two specialized computing engines for optimized performance. The AIMC engine handles feed-forward, embedding, and fully connected layers using spiking neuron tiles, with NVM crossbars for efficient weight storage. The SSA engine implements the MHSA mechanism in the spiking domain leveraging stochastic computing. This CMOS-based design enhances efficiency by replacing arithmetic units with logical AND gates.
@@ IV-A AIMC Engine
The AIMC engine is composed of multiple spiking neuron tiles, which are virtually grouped to ensure sufficient memory for storing weights in each feedforward or fully connected layer. Each tile comprises an input and output buffer, multiple synaptic arrays (SAs), and a group of LIF units. A shared DAC unit is integrated into each spiking neuron tile to facilitate the initial weight programming. This DAC unit is bypassed during inference, eliminating unnecessary overhead.
IV-A1 NVM Crossbars
Each SA includes a fixed-sized NVM crossbar array along with multiplexed readout units that allow MVMs to be performed in constant time, $O(1)$, directly within the memory, as introduced in Section II-D. This eliminates the need for energy-intensive CMOS arithmetic operations in conventional digital implementations, significantly reducing computational energy consumption. We impose a restriction on the size of crossbar arrays, because excessively large sizes result in increased interconnect resistance and capacitance, higher power consumption, signal degradation, and increased complexity of control circuitry. Consequently, the weights of a feedforward or fully connected layer are most likely distributed across multiple SAs or even several spiking neuron tiles.
We choose the phase change memory (PCM) device to implement the crossbars in SAs due to its multi-level cell capability (up to 4 bits per device), high endurance, fast read/write speed, and compatibility with CMOS technology [45, 46, 47]. Table II summarizes the detailed SA configuration parameters used in this work. Nevertheless, as an analog device, PCM is susceptible to thermal noise, $1/f$ noise, device variability, and quantization effects. A key challenge is conductance drift, where the stored conductance of a memory cell changes over time after programming. This drift can degrade neural network performance by affecting the precision and reliability of stored weights. Methods to address the effects caused by these nonidealities will be introduced in Section V.
TABLE II: Xpikeformer Configuration of Synaptic Arrays
@@ IV-A2 Weight Mapping and Tile Architecture
To efficiently manage this distribution and avoid the overhead of storing non-binary pre-activations, we adopt a row-block-wise mapping strategy. For example, with a crossbar size of 128$\times$128 cells and a weight matrix of 384$\times$512, the weight matrix must be divided into twelve 128$\times$128 submatrices. As shown in Fig. 4, the row-block-wise mapping strategy ensures that the same row of submatrices is mapped to the SAs within a single spiking neuron tile. For instance, rows 129-256 of the weight matrix are mapped to four SAs (2-1 to 2-4) within one spiking neuron tile. The input 512$\times$1 vector is then split into four 128$\times$1 subvectors and fed into SA 2-1 to SA 2-4, respectively.
At each SA column output of an SA, only a local sum of the vector dot product is obtained. To avoid storing this sum, the outputs from the same column of SA 2-1 to SA 2-4 are routed to a shared LIF unit, where they are accumulated using a carry-save adder (CSA) to produce a single pre-activation value. This pre-activation is then added to the current membrane potential via an adder.
The new membrane potential overwrites the previous one in a shift register. A comparator then compares this value against a threshold stored in the threshold register. If the membrane potential exceeds the threshold, the comparator generates a ‘1’ and resets the shift register. At each time step, the shift register undergoes a right shift by one bit, representing a leak factor of $\beta=0.5$. The comparator output, which acts as the spiking neuron’s output, is sent to the output buffer and subsequently stored in shared SRAM.
The row-block-wise mapping strategy is compatible with ADC sharing in SAs. As shown Fig. 4, the sharing ratio is 8, meaning each SA has 128///8$=$16 readout units. Consequently, SA 2-1 to SA 2-4 are connected to 16 LIF units. The address decoding strategy of the MUX decoder is identical across all SAs to ensure alignment between local sums. During each MUX decoding cycle, 16 out of 128 output features are obtained.
This row-block-wise mapping strategy minimizes data movement latency by directly routing local sums from multiple crossbars to LIF units, where they are accumulated and processed without intermediate buffering. This design reduces memory access bottlenecks, ensuring low latency and high energy efficiency.
Figure 4: A illustration of the row-block-wise mapping strategy.
@@ IV-B SSA Engine
Stochastic spiking attention (SSA) is a software-hardware co-designed spike-based attention mechanism optimized to minimize computing and memory access overhead. Unlike ANN-based attention, which relies on computationally intensive real-valued matrix multiplications, and SNN-based attention, which often uses masked integer addition, SSA utilizes stochastic computing to enable ultra-lightweight binary operations. This is made possible by the binary and stochastic nature of the queries $Q$, keys $K$, and values $V$, which are encoded using Bernoulli neurons. We introduce the SSA engine following a bottom-up approach.
@@ IV-B1 SSA Algorithm
where LIF($\cdot$) represents the per-entry application of LIF neuron as described in Section II-C2. The proposed SSA replaces the stateful LIF neurons with the Bernoulli neuron layer (BNL), computing the attention as:
@@
A detailed description of the SSA algorithm is shown in Algorithm 1. Owing to the binary nature of the representations $Q^{t}$, $K^{t}$, and $V^{t}$, the matrix multiplications in (6) are realized by lightweight logic AND operations and additions instead of full-precision multiplication units. In step 5, the $N\times N$ binary attention scores $S^{t}$ are obtained by first calculating the dot-product between $Q^{t}$ and $K^{t}$. This is done by accumulating the results of logical AND operations across the dimension $d_{K}$. The result is then normalized and used for Bernoulli encoding. Step 9 applies a similar logical AND to compute the matrix multiplication between $S^{t}$ and $V^{t}$ and generate probabilistically encoded attention result $A^{t}$.
@@
9:          $A^{t}_{d,n}\sim\mathrm{Bern}\left(\frac{1}{N}\sum_{n^{\prime}=1}^{N}S^{t}_{n,n^{\prime}}\land V^{t}_{d,n^{\prime}}\right)$
@@ IV-B2 SSA Tile
Each SSA tile is composed of $N^{2}$ physical stochastic attention cells (SACs) arranged in an $N\times N$ array. The $N^{2}$ parallelization level in SACs provides several significant benefits. First, it enables the simultaneous pairing of each query with every key-value pair. This is achieved by streaming $K^{t}$ and $V^{t}$ across columns and $Q^{t}$ across rows using 1-bit data buses. Additionally, this architecture allows for the parallel computation of all $N\times N$ elements of the attention score $S^{t}$ matrix. The $(i,j)$-th element of $S^{t}$ is computed within the corresponding $(i,j)$-th SAC. Given that SOTA transformers for edge AI applications typically operate with $N$ values ranging from 16 to 128, our design incorporates $N^{2}$ physical SAC units to ensure scalability and efficiency.
@@
To calculate the attention output $A^{t}$, the generated attention scores $S^{t}_{i,j}$ are held in the SAC for $d_{K}$ clock cycles. The local attention result between $S^{t}_{i,j}$ and the $j$-th row of $V^{t}$ is calculated using another AND gate, serving as the output of the SAC. To avoid using external delays for $V^{t}$, a $d_{K}$-bit shift register operating on a first-in-first-out basis is deployed in each SAC to temporarily buffer $V^{t}$ and align it with $S^{t}$. This allows for the simultaneous streaming of $Q^{t}$, $K^{t}$, and $V^{t}$ within the SSA block, facilitating pipelining over time steps. The summation of the local attention results is achieved by adding the outputs of each column of SACs using an $N$-input binary adder. The sum is then sent to another Bernoulli encoder, where it is normalized by $N$ and used as a probability to generate the Bernoulli samples. The $i$-th row of $A^{t}$ is generated sequentially by holding $S^{t}_{i,j}$ in the $(i,j)$-th SAU when streaming the $j$-th row of $V^{t}$. As a result of this procedure, the $i$th column within the SSA tile sequentially outputs the $i$-th row of the attention result $A^{t}$. With $N$ rows of SACs operating in parallel, the entire matrix $A^{t}$ is calculated column by column by one SSA tile.
@@
The Bernoulli encoder uses a comparator to compare the input integer with a pseudo-random number (PRN) generated from an external source. Notably, we do not perform normalization in hardware. Instead, we directly compare the unnormalized input integer $I$ with a random integer sampled uniformly from $(0,I_{\text{max}}]$. In our design, $I_{\text{max}}$ is set to $d_{K}$ within SACs and to $N$ outside of SACs. Choosing $d_{K}$ and $N$ as powers of two can streamline the hardware design by allowing directly using a fixed-point random integer sampled from a uniform distribution for the comparison. This approach simplifies the design and enables the use of power-efficient linear feedback shift register (LFSR)-based PRN generators.
@@ IV-B3 SSA Engine
The SSA engine is designed for efficient execution of the entire MHSA mechanism in spiking transformers. It consists of multiple SSA tiles, each serving as an individual attention head. These tiles are stateless, allowing them to be reused across different layers of the transformer, enhancing both area and energy efficiency. Additionally, the SSA engine incorporates an LFSR array that generates all the necessary PRNs for the Bernoulli neurons, ensuring a consistent supply of random numbers required for the stochastic elements of the model.
To further optimize power and area efficiency, the SSA engine implements a custom reuse strategy for random number generation, as described in [48]. Here we implement a 32-bit LFSR and tap all 4 bytes from the LFSR instead of just the lowest 8-bits [49]. This approach maximizes the utilization of generated random numbers, conserving both power and area. By leveraging these design strategies, the SSA engine achieves a balance of high performance and resource efficiency.